MINUTES FROM THE BOOSTER SUPERVISOR MEETING
Links: Minutes, PSB home page, PSB
PRESENT: M. Benedikt, A. Blas, M. Chanel, G. Cyvoct, E.
Jensen, M. Lindroos, K. Schindl, R. Scrivens.
Supervisor Linac: M. O'Neil à C. Dutriat
Supervisor PSB: E. Jensen à M. Benedikt
- Linac 2: Week without any major problems.
- The reset procedure for Linac2 buncher remote control is available on the
web via the PSB page or directly under http:
- Linac 3: Back to lead ion operation since last Wednesday. The beam
is used by E. Mahner for LEIR vacuum chamber tests.
- Wednesday: BTM/BTP pick-ups were perturbed by an offset that disappeared
after changing a NIM crate power supply.
- Saturday: there were again some problems with the active filter for the
ejection septum BE.SMH15L1.
- Monday: Ejection kicker ring1, BE1.KFA14L1 was down with hardware
interlock. A power supply card needed to be changed (L.Sermeus).
- Please note that all the RF GFAS are now programmed in C-time and no
longer with internal stops. The new start timing for all the GFAS is
BAX.SGSRF in the working set RF global control.
- Status of the different beams:
- The 2nd harmonics component at extraction was reduced to the minimum
value that just compensates for the beam induced voltage (~0.4 kV).
Needed to ensure correct longitudinal matching with the PS.
- Double batch injection in the PS (LHC/TSTLHC in PSB) works fine.
- Setup started last Tuesday, some more work needed for the bunch
- Too big transverse emittances were found on Monday afternoon.
Readjustment of injection, Q-strips etc. resolved the problem.
- During staggered operation (only rings 2,3,4) a watchdog problem was
encountered. To resolve the problem the start timing and the kick of the
extraction kicker ring 1 need to be disabled (BEX.SKFA14L1 disabled,
BE.KFA14L1 ->kick off).
High intensity beams:
- The capture process was optimized on all high intensity beams last
- Next BS meeting Tuesday 3-June-02 at 10.00 in the MCR Glassbox.
Minutes by M. Benedikt.